OpenStudy (anonymous):

the mathematical value of 0! is what?

OpenStudy (anonymous):

1

terenzreignz (terenzreignz):

It's 1, isn't it? Due the fact that \[\Large n! = n(n-1)!\] And just take n = 1

terenzreignz (terenzreignz):

You get \[\Large 1! = 1(1-1)! = 1\cdot 0! \]

OpenStudy (anonymous):

yeah I thought it equaled 1 but wasn't sure

terenzreignz (terenzreignz):

It does... either by convention, or something else ^_^

OpenStudy (anonymous):

cool, thanks